# Small-scale magnetohydrodynamic dynamos: from deterministic chaos to turbulence

### v3

@@ II.2 Distributed chaos in magnetic field driven by magnetic helicity
One can find the probability distribution $P(k_{c})$ for the fluid (or plasma) dynamics dominated by the magnetic helicity using the dimensional considerations and a scaling relationship
@@
relating the characteristic value of the magnetic field $B_{c}$ to the characteristic value of the wavenumber $k_{c}$.
If the characteristic value of the magnetic field $B_{c}$ has a normal distribution (with zero mean my ), then it follows from Eq. (7) that the characteristic value of the wavenumber $k_{c}$ has the chi-squared ($\chi^{2}$) distribution
@@ III Magneto-inertial range of scales
A relationship between the exponents $\beta$ and $\alpha$ can be readily obtained (using some algebra) from the Eqs. (17) and (19) for the normally distributed $B_{c}$
@@ VI Geodynamo
It is now believed that the magnetic field of Earth is generated by a flow of liquid metal in the Earth’s core (the geodynamo). The measurements inside the core are impossible while the systematic good quality measurements of the spatial distribution of the magnetic field on the Earth’s surface are still rare. Therefore, direct numerical simulations can be rather useful. The problems with the DNS of the geodynamo are also numerous. First of all, it is still very difficult to approach the values of the parameters characteristic for the real geodynamo (see, for instance, Table 1 in a recent paper Ref. sjna ).
In the Ref. sjna a recent attempt to this end was made using maximal possibilities of the present computers. The authors of the DNS numerically solved the equations of the thermally (buoyancy) driven convection of an electrically conducting fluid with a chemical species concentration gradient (in the Boussinesq approximation with the Lorentz force) in a rotating spherical shell mimicking the liquid geodynamo confined between the mantle and solid inner core:
@@
The time-averaged spatial spectral data at the spheres of different radii $R$ were provided in the paper Ref. sjna for the case with the most vigorous motion: $Re=5140$, $Ra=2.54\times 10^{12}$, $Ro=5.1\times 10^{-4}$, $E=10^{-7}$, $Pr=1$, $Re_{m}=514$, and $Pr_{m}=0.1$. The system is helical, and for this case the bulk relative helicity and helicity at the boundaries were also rather strong.
It should be noted that the spectral data were provided in Ref. sjna as a dependence of the energy density spectra $E(l)$ on the spherical harmonic degree $l$. On a sphere of radius $R$ the wavenumber $k$ is related to $l$ as
$kR=\sqrt{l(l+1)}$ (42)
@@
Figure 14 shows the spectral data taken from Fig. 2 of a recent paper finlay ). The spectral computations were produced using the CHAOS-7 model of the near-Earth magnetic field for the period 1999-2020 based on the observations made by the low-Earth orbit satellites CryoSat-2, Swarm, CHAMP, SAC-C, Orsted, and SAC-C, and on the ground observatory measurements. The dashed curve indicates the helical spectrum Eq. (9) (here $R$ is the Earth’s mean radius).

### v4

@@ II.2 Distributed chaos in magnetic field driven by magnetic helicity
It is well-known that in a weakly non-ideal case magnetic helicity is still almost conserved (since there are no processes that can effectively transfer it to the resistive scales) while the magnetic and total energy are efficiently dissipated, i.e. magnetic helicity is an adiabatic invariant in weakly non-ideal MHD. Therefore one can find the probability distribution $P(k_{c})$ for the fluid (or plasma) dynamics dominated by the magnetic helicity using the dimensional considerations and a scaling relationship
@@
relating the characteristic value of the magnetic field $B_{c}$ to the characteristic value of the wavenumber $k_{c}$ through magnetic helicty $|h_{m}|$ (as an adiabatic invariant).
If the positive variable $B_{c}$ has the half-normal probability distribution $P(B_{c})\propto\exp-(B_{c}^{2}/2\sigma^{2})$ (my, ) (i.e. a normal distribution with zero mean truncated to only have nonzero probability density for positive values of its argument: if $B$ is a normal random variable, then $B_{c}=|B|$ has a half-normal distribution jkb ), then it follows from Eq. (7) that the characteristic value of the wavenumber $k_{c}$ has the chi-squared ($\chi^{2}$) distribution
@@ III Magneto-inertial range of scales
A relationship between the exponents $\beta$ and $\alpha$ can be readily obtained (using some algebra) from the Eqs. (17) and (19) for the half-normally distributed positive variable $B_{c}$
@@ VI Geodynamo
Figure 10: Magnetic energy spectrum $E(k)$ generated by a chaotic magnetohydrodynamic dynamo in a conducting spherical Couette flow at $Pr_{m}=0.25,1,4$ (a direct numerical simulation).
It is now believed that the magnetic field of Earth is generated by a flow of liquid metal in the Earth’s core (the geodynamo). The measurements inside the core are impossible while the systematic good quality measurements of the spatial distribution of the magnetic field on the Earth’s surface are still rare. Therefore, direct numerical simulations can be rather useful.
Figure 10 shows the magnetic energy spectra $E(k)$ (against the wavenumber $k$) generated by a magnetohydrodynamic dynamo in a conducting spherical Couette flow at the magnetic Prandtl numbers $Pr_{m}=\nu/\eta=0.25,1,4$ (here $\nu$ is the kinematic viscosity and $\eta$ is the magnetic diffusivity of the fluid) and the Ekman number $E=\nu/(\Omega r_{o}^{2})=10^{-3}$ (here $\Omega$ is the rotational rate of the outer sphere and $r_{o}$ is its radius). This is a flow between two concentric spheres with a differential rotation relevant to the Earth’s outer liquid core. To simulate the situation in the Earth’s core the aspect ratio $r_{i}/r_{o}=0.35$ (here $r_{i}$ is the radius of the inner core, see for more details gc ).
The spectral data were taken from Fig. 13 of the Ref. (gc, ). The spectra were time and $r$ averaged over the whole shell. $R$ is an effective radius ($r_{i}<R<r_{o}$). It should be noted that the spectral data were provided in (gc, ) as a dependence of the energy density spectra $E(l)$ on the spherical harmonic degree $l$. On a sphere of radius $R$ the wavenumber $k$ is related to $l$ as
$kR=\sqrt{l(l+1)}$ (38)
The dashed curves indicate the best fit with the exponential Eq. (2) (the deterministic chaos), and stretched exponentials Eqs. (32) and (34) (the dissipative distributed chaos). Randomization increases with the $Pr_{m}$ in this case. The dotted vertical arrow indicates the position of the $k_{c}$. As one can see the small-scale dynamo coexists with the large-scale dynamo and the large-scale coherent structures determine the deterministic chaos (the position of the $k_{c}R$ coincides with the position of the spectral peak indicating the coherent structures).
It is still very difficult to approach the values of the parameters characteristic for the real geodynamo (see, for instance, Table 1 in a recent paper Ref. sjna ). In the Ref. sjna a recent attempt to this end was made using maximal possibilities of the present computers. The authors of the DNS numerically solved the equations of the thermally (buoyancy) driven convection of an electrically conducting fluid with a chemical species concentration gradient (in the Boussinesq approximation with the Lorentz force) in a rotating spherical shell mimicking the liquid geodynamo confined between the mantle and solid inner core:
@@
The difference in the ocean and land spectra could be related to the electric conductivity of the mantle (see, for instance, shebalin2 and references therein) which can be different under ocean and land. This provides different local boundary conditions for the outer liquid core below the ocean and the land.
@@
Figure 15 shows (in the semi-log scales) the spectral data taken from Fig. 2 of a recent paper finlay ). The spectral computations were produced using the CHAOS-7 model of the near-Earth magnetic field for the period 1999-2020 based on the observations made by the low-Earth orbit satellites CryoSat-2, Swarm, CHAMP, SAC-C, Orsted, and SAC-C, and on the ground observatory measurements. The dashed curve indicates the helical spectrum Eq. (9) (here $R$ is the Earth’s mean radius).